the larger of two positive integers is five more than twice the smaller integer. The product of the integers is 52. What are the integers?
Answer by checkley77(12569)

y=2x+5
xy=52
x(2x+5)=52
2x^2+5x-52=0
(2x+13)(x-4)=0
x-4=0
x=4 ans.
y=2*4+5
y=8+5
y=13 ans.
Proof:
4*13=52
52=52